Need the solution of this problem . Thanks !
A car is moving along a street with acceleration
Answer #1

The car distance is given by:

d=vot+at^2/2

but the distance of the train is also variable

dt=200+v't

making both equations equal:

200+v't=vot+at^2/2

200+10t=17t+0.5t^2/2

rearraging the expression

0=0.25t^2+7t-200

solving for t

t=17.559 s

we will obtain a - and a + solution, we take the + one